A beam of electrons passes undeflected through mutually perpendicular electric and magnetic fields of magnitudes 6 kVm^-1 and 10^-3 T, respectively. By maintaining the same magnetic field when the electric field is cut off, the electrons move in the magnetic field in a circular path. Calculate the speed of electrons and the radius of the circular path.
1
Expert's answer
2019-05-21T08:47:18-0400

The absence of deflection in the first part of the problem means that both electric and magnetic forces act along the same line but in the opposite directions, i.e.


"eE=evB \, \Rightarrow \, v = \frac{E}{B}"

Substituting the numerical values, we obtain


"v = \frac{6 \cdot 10^3}{10^{-3}} = 6 \cdot 10^6 \, m/s"

In case when only the magnetic field is maintained, the condition of the circular orbit states:


"evB = \frac{m v^2}{R} \, \Rightarrow \, R = \frac{m v}{eB} = \frac{m E}{e B^2}"

Substituting the numerical values, we obtain


"R = \frac{9.1 \cdot 10^{-31} \cdot 6 \cdot 10^{3}}{1.6 \cdot 10^{-19} \cdot 10^{-6}} \approx 3.4 \cdot 10^{-2} \, m = 3.4 \, cm"

Answer: 6x106 m/s; 3.4 cm